OpenStudy (poetryprincess):

Could some one please explain to me how to subtract positive and negative numbers? For example if there was like (-95)-(-53)= ?

OpenStudy (jiteshmeghwal9):

\[\Large -*-=+\]

OpenStudy (jiteshmeghwal9):

-95-(-53)=-95+53=-(95-53)

OpenStudy (poetryprincess):

Two negatives are suppose to equal a negative.

OpenStudy (poetryprincess):

I figured it out. Thanks anyway. It was -95+53=-42